OpenStudy (anonymous):

Evaluate the expression for the given value of the variable(s). -x^2 - 4x; x = –3 A. –17 B. 11 C. –1 D. 3

OpenStudy (ddcamp):

-(-3)² - 4(-3) = -(9) - (-12) = -9 + 12 = 3, D

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

is it right? @ganeshie8

ganeshie8 (ganeshie8):

its perfect !

ganeshie8 (ganeshie8):

you should see it by yourself, the working... :)

OpenStudy (anonymous):

ok